# Market-Driven Equilibria for Distributed Solar Panel Investment

### v1

@@ Abstract
This study investigates market-driven long-term investment decisions in distributed solar panels by individual investors. We consider a setting where investment decisions are driven by expected revenue from participating in short-term electricity markets over the panel’s lifespan. These revenues depend on short-term markets equilibria, i.e., prices and allocations, which are influenced by aggregate invested panel capacity participating in the markets. We model the interactions among investors by a non-atomic game and develop a framework that links short-term markets equilibria to the resulting long-term investment equilibrium. Then, within this framework, we analyze three market mechanisms: (a) a single-product real-time energy market, (b) a product-differentiated real-time energy market that treats solar energy and grid energy as different products, and (c) a contract-based panel market that trades claims/rights to the production of certain panel capacity ex-ante, rather than the realized solar production ex-post. For each, we derive expressions for short-term equilibria and the associated expected revenues, and analytically characterize the corresponding long-term Nash equilibrium aggregate capacity. We compare the solutions of these characterizing equations under different conditions and theoretically establish that the product-differentiated market always supports socially optimal investment, while the single-product market consistently results in under-investment. We also establish that the contract-based market leads to over-investment when the extra valuations of users for solar energy are small. Finally, we validate our theoretical findings through numerical experiments.
@@ I Introduction
The increasing adoption of renewable energy, particularly solar photovoltaic systems, plays an important role in modern power systems. While utility-scale solar projects contribute significantly to overall capacity, distributed rooftop installations are playing an increasingly prominent role in reshaping the electricity landscape. Notably, as of 2022, approximately 140 GW of solar capacity had been installed in the U.S., with more than 35% of it coming from installations by households and commercial entities [1].
Unlike centralized projects planned and operated by utilities, distributed solar investments are driven by decisions of individual households and businesses. Several factors influence these decisions. A key factor is the financial return from the investment, driven by the expected revenue earned in short-term distribution-level markets over the panels’ lifespan, where that revenue is determined by the markets’ equilibria. Meanwhile, these short-term equilibria depend on the aggregate invested panel capacity available in the markets, which is shaped by the long-term investment equilibrium that emerges from interactions among individual investors. Therefore, there is a feedback loop between short-term markets equilibria and long-term investment equilibrium, as each influences the other. A precise characterization of this relationship is crucial for understanding the emergence of investment equilibria. This insight is not only important for utilities and independent system operators, who should account for distributed capacity growth in their planning studies, but also for policymakers and market designers aiming to understand the long-term impact of diverse market designs on the solar panel adoption. This motivates us to study the following critical open question: How do short-term market equilibria, arising from diverse solar market mechanisms, shape the corresponding long-term distributed investment equilibrium?
@@ I-A Organization and Contributions
To address this problem, we consider a large population of individual investors and aim to characterize the long-term investment equilibrium that emerges from the previously described feedback loop. We first develop an investment model, formulated as a non-atomic game, that links long-term investment decisions to expected revenues under a given market design (Section II). To characterize these revenues, we introduce a unified modeling framework for short-term markets (Section III). This framework allows us to formally define and compare different market designs by capturing key features such as product differentiation and ex-ante/ex-post trading. We then apply it to three representative mechanisms that differ in these key features (Section IV): (a) a single-product real-time market, (b) a product-differentiated real-time market, and (c) a contract-based panel market. For each, we first derive analytical characterizations of the short-term market equilibria and then use the investment model to derive the associated long-term equilibrium aggregate capacities (Section V). Our analysis, initially assuming homogeneous operation periods, is extended to handle heterogeneous operation periods (Section VI), followed by numerical experiments (Section VII).
@@
    The non-atomic game-based investment model encapsulates how individual investors impact each other through their aggregate investment decisions, offering an analytically tractable way to link short-term market equilibria to long-term aggregate investment equilibrium.
@@
    The proposed unified modeling structure provides closed-form characterizations of short-term market equilibria across diverse mechanism designs, accommodating both ex-ante and ex-post trading as well as heterogeneous consumer preferences.
@@
    These analytical characterizations enable us to theoretically establish that the single-product real-time market, where solar energy is pooled with conventional grid energy and traded ex-post, leads to under-investment. In contrast, the product-differentiated market, where solar energy is traded ex-post and may be sold at a premium due to heterogeneous user preferences, yields the socially optimal capacity. Finally, the contract-based market, where ex-ante panel capacity is traded instead of realized generation, can result in over-investment when users’ extra valuations for solar are low.
@@ I-B Related Literature
This work is related to the intersection of two broad lines of literature: distribution-level electricity market mechanisms and long-term investment in solar panels. Most existing studies on the investment side of the literature adopt a centralized perspective, focusing on the optimal strategy of a single decision-maker such as a utility planner or an individual investor (see, e.g., [2, 3, 4, 5, 6]). However, this line of work neglects the interdependencies of investment decisions in distributed settings, where the actions of one investor can influence the returns of others through shared market mechanisms.
On the market side, a variety of distribution-level electricity market designs have been proposed. Understanding how these structures differ is essential for our analysis, as their design features shape the expected revenue of investors. In this context, prior research has proposed a range of market designs, with particular emphasis on real-time pricing and service contract-based approaches. In a broad sense, real-time pricing mechanisms involve the determination of time-varying electricity prices based on the ex-post realized values of supply and demand. In contrast, service contract-based mechanisms typically offer less volatile prices in exchange for a fixed level of electricity service, relying on ex-ante statistics such as those related to random solar generation. Extensive research has explored the relative merits of these retail market designs, examining factors such as economic efficiency [7, 8, 9, 10], price and bill volatility [11, 12, 13], and environmental impacts [14]. In addition, some studies have examined another facet of market design: whether electricity from solar and conventional sources should be treated as differentiated products in electricity markets, given their distinct operational characteristics and emissions [15, 16, 17, 18, 19]. Despite the contributions of all these reviewed studies, they all focus primarily on short-term performance criteria and overlook the long-term impacts of their designed markets on investment decisions.
The most closely related paper to ours is [20], which analyzes the equilibrium panel investment capacities under feed-in-tariff and sharing economy markets. Nonetheless, they adopt a finite-player game to analyze the setup involving large investors. In contrast, we focus on the setting with a large number of small investors. Through a non-atomic game formulation, we can analytically compare equilibrium capacities driven by various market mechanisms, for which they have to resort to numerical simulation.
@@ II Investment Model
A large group of potential investors considers installing solar panels and becoming panel owners. Their decisions weigh upfront capital and installation costs against expected revenues from participating in short-term electricity markets over the panels’ lifespan. An overview of the setup is shown in Fig. 1.
@@ II-A Connecting Planning and Operation Timescales
Investors typically consider a 25-year lifespan for solar panels [22]. However, estimating expected revenue at the time of investment is challenging, as reliable long-term forecasts of solar output and market conditions are rarely available. A common approach is to use a shorter planning period (e.g., one year), estimate revenues from historical data, and scale [^1] This step may include interest rate adjustments and net present value computations. the result by the number of such intervals over the panel’s lifespan. The appropriate length of the planning period should be selected based on the availability and granularity of historical data.
When the planning period is long, there may be multiple operation periods within it. We consider a finite-horizon discrete-time planning window consisting of operation periods $\mathcal{T}=\{1,\dots,T\}$, indexed by $t$, where $T$ is the total number of such periods. For each operation period $t\in\mathcal{T}$, we model the short-term operation conditions (e.g., supply and demand) as constants and incorporate variability (e.g., of solar generation) across different operation periods by sampling from a probability distribution. Thus, the expected revenue remains consistent across operation periods. We estimate it for a single representative period (e.g., as shown in Fig. 1), then scale it up to the planning horizon and ultimately to the full lifespan. The total expected revenue is therefore obtained by multiplying the representative period’s revenue by a scaling factor $\widetilde{T}$, which captures the number of such periods over the panel’s lifespan and incorporates any interest-rate adjustments.
@@ II-B Panel Investment Game
Considering a large number of small investors each of whom does not have market power, we model the group of investors as the continuous interval $\mathcal{I}_{\mathrm{inv}}:=[0,1]$. We consider the case where each investor $i\in\mathcal{I}_{\mathrm{inv}}$ decides whether to install a solar panel with a fixed capacity $\overline{c}$. Let the decision of investor $i$ be $x_{i}\in\{0,1\}$. The total panel capacity that the investors will install is then
$c=\overline{c}\int_{\mathcal{I}_{\mathrm{inv}}}x_{i}\,\,\mathrm{d}i.$ (1)
@@
${\color[rgb]{0,0,0}c}\pi_{0}\leq\widetilde{T}\pi_{\mathrm{u}}\mathbb{E}[{\color[rgb]{0,0,0}c}G],$ (2)
where $\pi_{\mathrm{u}}>0$ is the fixed price at which the utility provides backstop electricity to buyers, and $G$ denotes solar generation per unit panel capacity, typically measured in $\mathrm{W}/\mathrm{m}^{2}$ or $\mathrm{GW}/\mathrm{km}^{2}$. This solar generation is modeled as a random variable with a cumulative distribution function $F_{G}$ and a probability density function $f_{G}$. For simplicity, we assume that $f_{G}$ is differentiable and that $G$ is supported on $\mathbb{R}_{+}$ with finite first and second moments, and $f_{G}(g)>0$ for all $g\in\mathbb{R}_{+}$.
@@
$\Pi^{\mathrm{inv}}_{m,i}(x_{i},c)=\left[(\Pi_{m}^{\mathrm{s}}(c)/c)-\pi_{0}\right]\overline{c}\,x_{i},$ (3)
where $\Pi_{m}^{\mathrm{s}}(c)$ represents the expected total revenue received by solar panel owners under the short-term market $m$.
Given the players $\mathcal{I}_{\mathrm{inv}}$, actions $x=\{x_{i}\}_{i\in\mathcal{I}_{\mathrm{inv}}}$, and payoff (3), we have defined the solar panel investment game, which is an aggregate game where the dependence of investor $i$’s payoff on the other investors’ decisions is only through the total panel capacity $c$. A Nash equilibrium (NE) for this game is defined as:
@@
For a market mechanism $m$, the aggregate panel capacity $c$ constitutes an NE for the solar panel investment game with market structure $m$ if there exists a collection of investment decisions $x$ such that $\Pi_{m,i}^{\mathrm{inv}}(x_{i},c)\geq\Pi_{m,i}^{\mathrm{inv}}(x_{i}^{\prime},c)$, for all $x_{i}^{\prime}\in\{0,1\}$ and $i\in\mathcal{I}_{\mathrm{inv}}$, and (1) holds.
@@ III-A Common Model Elements
Consider an operation period (e.g., the representative operation period in Fig. 1) where the expected revenue of investors is shaped by short-term conditions in the distribution-level solar energy market. The market consists of buyers (i.e., electricity consumers), sellers [^2] We acknowledge the possibility that electricity prosumers can be both buyers and sellers at the same time. For the market mechanisms that we will consider in this paper, a prosumer can be modeled as the superposition of a buyer and a seller. This observation may not apply to other markets where buying and selling prices are non-uniform. (i.e., panel owners), and a utility company that provides backstop supply when solar generation alone cannot fully satisfy demand. Similar to the investment game, we model the groups of a large number of buyers and sellers as continuous intervals.
Define the set of buyers by the interval $\mathcal{I}_{\mathrm{b}}:=[0,1]$. In our analysis, we consider a situation where environmentally conscious electricity consumers have an extra heterogeneous valuation for solar energy over the conventional energy supplied by the utility. For a buyer $i\in\mathcal{I}_{\mathrm{b}}$, we denote their extra valuation for each unit of load served by solar as $v_{i}\in\mathbb{R}_{+}$, referred to as the solar premium. The distribution of $v_{i}$ can be characterized by a function $F_{V}$ defined as $F_{V}(v)=\int_{i\in\mathcal{I}_{\mathrm{b}}}\,\mathbb{1}\{v_{i}\leq v\}\,\,\mathrm{d}i$, where $\mathbb{1}$ is the indicator function. For simplicity, we assume that $F_{V}$ is differentiable and strictly increasing on $0,\overline{v}]$, where $\overline{v}\geq 0$. We denote its derivative by $f_{V}$, supported on $0,\overline{v}]$, and assume $f_{V}(v)$ is bounded away from zero for all $v\in[0,\overline{v}]$. Let the total electric load of all consumers be $L$. We assume $L$ is inelastic, deterministic, and equally distributed across all consumers.

### v2

# Extended Version: Characterizing Distributed Photovoltaic Panel Investment Equilibria

@@ Abstract
This study investigates long-term investment decisions in distributed photovoltaic panels by individual investors. We consider a setting where investment decisions are driven by expected revenue from participating in short-term electricity markets over the panel lifespan. These revenues depend on short-term markets equilibria, i.e., prices and allocations, which are influenced by aggregate invested panel capacity participating in the markets. We model the interactions among investors by a non-atomic game and develop a framework that links short-term markets equilibria to the resulting long-term investment equilibrium. Then, within this framework, we analyze three market mechanisms: (a) a single-product real-time energy market, (b) a product-differentiated real-time energy market that treats solar energy and grid energy as different products, and (c) a contract-based panel market that trades claims/rights to the production of certain panel capacity ex-ante, rather than the realized solar production ex-post. For each, we derive expressions for short-term equilibria and the associated expected revenues, and analytically characterize the corresponding long-term Nash equilibrium aggregate capacity. We compare the solutions of these characterizing equations under different conditions and theoretically establish that the product-differentiated market always supports socially optimal investment, while the single-product market consistently results in under-investment. We also establish that the contract-based market leads to over-investment when the extra valuations of users for solar energy are small. Finally, we validate our theoretical results through numerical experiments.
@@ I Introduction
The increasing adoption of renewable energy, particularly photovoltaic (PV) panels, plays an important role in modern power systems. While utility-scale solar projects contribute significantly to overall capacity, distributed rooftop installations are playing an increasingly prominent role in reshaping the electricity landscape. Notably, by the end of 2024, the U.S. had installed about 175 GW of solar capacity, over 31% of which came from residential and commercial installations [1].
Unlike centralized projects planned and operated by utilities, distributed solar investments at the distribution level reflect individual adoption and investment decisions by households and businesses. Several factors influence these decisions. A key factor is the financial return from the investment, driven by the expected revenue earned in short-term distribution-level markets over the panel lifespan, where that revenue is determined by the markets’ equilibria. Meanwhile, these short-term equilibria depend on the aggregate invested panel capacity available in the markets, which is shaped by the long-term investment equilibrium that emerges from interactions among individual investors. Therefore, there is a feedback loop between short-term markets equilibria and long-term investment equilibrium, as each influences the other. A precise characterization of this relationship is crucial for understanding the emergence of investment equilibria. This insight is not only important for utilities and independent system operators, who should account for distributed capacity growth in their planning studies, but also for policymakers and market designers aiming to understand the long-term impact of diverse market designs on PV panel adoption. This motivates us to study the following critical open question [^1] We note that a few studies in related settings (e.g., conventional generation investment) also link long-run investment equilibria to short-run operational equilibria. However, they largely focus on spot markets [2, 3] and either do not yield closed-form characterizations [2] or incorporate detailed consumer modeling (e.g., heterogeneity) [2, 3].: How do short-term market equilibria, arising from diverse solar market mechanisms, shape the corresponding long-term distributed investment equilibrium?
@@ I-A Organization and Contributions
To address this open problem, we consider a large population of individual investors and aim to characterize the long-term investment equilibrium that emerges from the previously described feedback loop. We first develop an investment model, formulated as a non-atomic game, where each player is infinitesimal with negligible market power, reflecting the large number of small investors in distribution-level settings. This model links long-term investment decisions to expected revenues under a given market design (Section II). To characterize these revenues, we introduce a unified modeling framework for different types of short-term markets (Section III) that allows us to formally define and compare different market designs. Our unified market model captures key features such as product differentiation and ex-ante/ex-post trading. We then apply it to three representative mechanisms that differ in these features (Section IV). First, we study (a) a single-product real-time market similar to today’s real-time electricity markets; then (b) a new product-differentiated real-time market to assess the impact of distinguishing clean solar energy from conventional grid energy; and finally (c) a new contract-based panel market that operates on a much slower timescale than real-time markets. For each, we first derive analytical characterizations of the short-term market equilibria and then use the investment model to obtain and compare the associated long-term equilibrium aggregate capacities (Section V). Our analysis, initially assuming homogeneous operation periods, is extended to handle heterogeneous operation periods (Section VI), followed by numerical experiments (Section VII).
@@
    We offer an analytically tractable way to link short-term market equilibria to long-term aggregate investment equilibrium through a non-atomic game-based investment model that captures how individual investors affect one another through aggregate investment decisions. This model can also represent common pricing schemes (e.g., feed-in tariffs and time-of-use rates) as long-run averages of short-term market prices that co-evolve with investment decisions, rather than exogenous fixed tariffs.
@@
    We develop a unified market-modeling framework that delivers closed-form characterizations of short-term market equilibria across diverse market mechanisms, accommodating both ex-ante and ex-post trading as well as heterogeneous consumer preferences.
@@
    These analytical characterizations yield new insights into how different market mechanisms shape PV investment. Specifically, we theoretically establish that the single-product real-time market, where solar energy is pooled with conventional grid energy and traded ex-post, leads to under-investment. Thus, we believe there is significant value in considering alternative market mechanisms. In particular, the product-differentiated market, where solar energy is traded ex-post and may be sold at a premium due to heterogeneous user preferences, yields the socially optimal capacity. Finally, the contract-based market, where ex-ante panel capacity is traded instead of realized generation, can result in over-investment when users’ extra valuations for solar energy are small. Such over-investment can be desirable under economies of scale, where higher aggregate deployment lowers the cost of PV investment.
@@ I-B Related Literature
This work is related to the intersection of two broad lines of literature: distribution-level electricity market mechanisms and long-term investment in solar panels. Most existing studies on the investment side of the literature adopt a centralized perspective, focusing on the optimal strategy of a single decision-maker such as a utility planner or an individual investor (see, e.g., [4, 5, 6]). However, this line of work neglects the interdependencies of investment decisions in distributed settings, where the actions of one investor can influence the returns of others through shared market mechanisms.
On the market side, a variety of distribution-level electricity market designs have been proposed. Understanding how these structures differ is essential for our analysis, as their design features shape the expected revenue of investors. In this context, prior research has proposed a range of market designs, with particular emphasis on real-time pricing and service contract-based approaches. In a broad sense, real-time pricing mechanisms involve the determination of time-varying electricity prices based on the ex-post realized values of supply and demand. In contrast, service contract-based mechanisms typically offer less volatile prices in exchange for a fixed level of electricity service, relying on ex-ante statistics such as those related to random solar generation. Extensive research has explored the relative merits of these retail market designs, examining factors such as economic efficiency [7, 8], price and bill volatility [9, 10], and environmental impacts [11]. In addition, some studies have examined another facet of market design: whether electricity from solar and conventional sources should be treated as differentiated products in electricity markets, given their distinct operational characteristics and emissions [12, 13, 14]. Despite the contributions of these studies[7, 8, 9, 10, 11, 12, 13, 14], they all focus primarily on short-term performance criteria and overlook the long-term impacts of their designed markets on investment decisions.
The most closely related paper to ours is [15], which studies equilibrium PV investment capacities under feed-in-tariff and sharing-economy markets. However, it adopts a finite-player game with large investors, whereas we consider a large population of small investors. Our non-atomic game formulation enables analytical comparisons of equilibrium capacities driven by different market mechanisms, for which they have to resort to numerical simulations. While non-atomic game formulations have been used in other contexts (e.g., workplace EV charging pricing [16] and distributed storage investment [17]), they have not been used in an investment framework that links short-term market equilibria under diverse market designs to long-term investment equilibrium.
@@ II Investment Model
A large group of potential investors considers installing PV panels and becoming panel owners. Their decisions weigh upfront capital and installation costs against expected revenues from participating in short-term electricity markets over the panel lifespan. An overview of the setup is shown in Fig. 1.
@@ II-A Connecting Planning and Operation Timescales
Investors typically assume a 25-year PV panel lifespan [19]. Accordingly, at the investment stage, estimating expected revenue over such a long horizon is challenging because long-term forecasts of solar generation and market conditions are rarely available. In response, a common approach is to use a shorter planning period (e.g., one year), estimate revenues from historical data, and scale [^2] This step may include interest rate adjustments and net present value computations. the result by the number of such intervals over the panel lifespan. The appropriate length of the planning period should be selected based on the availability and granularity of historical data.
Since the planning period is long, it contains multiple operation periods, each on a shorter timescale (e.g., an hour or a day). We consider a finite-horizon discrete-time planning window consisting of operation periods $\mathcal{T}=\{1,\dots,T\}$, indexed by $t$, where $T$ is the total number of such periods. For each operation period $t\in\mathcal{T}$, we model the short-term operation conditions (e.g., supply and demand) as constants and incorporate variability (e.g., of solar generation) across different operation periods by sampling from a probability distribution. Thus, the expected revenue remains consistent across operation periods. Therefore, we can first estimate this revenue over a single operation period, which we refer to as the representative operation period as shown in Fig. 1, and then scale it to obtain the expected revenue for the planning horizon and, ultimately, for the panel lifespan. The total expected revenue is therefore obtained by multiplying the representative operation period’s revenue by a scaling factor $\widetilde{T}$, which captures the number of all operation periods over the panel lifespan and incorporates any interest-rate adjustments.
@@ II-B Panel Investment Game
Considering a large number of small investors, each of whom does not have market power and is infinitesimal relative to the market (so individual actions do not affect outcomes while the aggregate does), we model their competition as a non-atomic game [20]. Consequently, we represent the investors population as the continuum $\mathcal{I}_{\mathrm{inv}}:=[0,1]$. We consider the case where each investor $i\in\mathcal{I}_{\mathrm{inv}}$ decides whether to install a PV panel with a fixed capacity $\overline{c}$. Let the decision of investor $i$ be $x_{i}\in\{0,1\}$. The total panel capacity that the investors will install is then
$c=\overline{c}\int_{\mathcal{I}_{\mathrm{inv}}}x_{i}\,\,\mathrm{d}i.\vskip-2.5pt$ (1)
@@
${\color[rgb]{0,0,0}c}\pi_{0}\leq\widetilde{T}\pi_{\mathrm{u}}\mathbb{E}[{\color[rgb]{0,0,0}c}G],\vskip-1.99997pt$ (2)
where $\pi_{\mathrm{u}}>0$ is the fixed price at which the utility provides electricity to buyers, and $G$ denotes solar generation per unit panel capacity, typically measured in $\mathrm{W}/\mathrm{m}^{2}$ or $\mathrm{GW}/\mathrm{km}^{2}$. This solar generation is modeled as a random variable with a cumulative distribution function $F_{G}$ and a probability density function $f_{G}$. For simplicity, we assume that $f_{G}$ is differentiable and that $G$ is supported on $\mathbb{R}_{+}$ with finite first and second moments, and $f_{G}(g)>0$ for all $g\in\mathbb{R}_{+}$.
@@
$\Pi^{\mathrm{inv}}_{m,i}(x_{i},c)=\left[(\Pi_{m}^{\mathrm{s}}(c)/c)-\pi_{0}\right]\overline{c}\,x_{i}.$ (3)
Here, $\Pi_{m}^{\mathrm{s}}(c)$ represents the expected total revenue received by panel owners under the short-term market $m$, to be defined individually for each market in Section III. Each investor $i\in\mathcal{I}_{\mathrm{inv}}$ chooses its investment decision to maximize $\Pi^{\mathrm{inv}}_{m,i}(x_{i},c)$.
Given the players $\mathcal{I}_{\mathrm{inv}}$, actions $x=\{x_{i}\}_{i\in\mathcal{I}_{\mathrm{inv}}}$, and payoff (3), we have defined the panel investment game, which is an aggregate game where investor $i$’s payoff depends on the other investors’ decisions only through the total panel capacity $c$. A Nash equilibrium (NE) for this game is defined as:
@@
The aggregate panel capacity $c$ constitutes an NE for the panel investment game under market mechanism $m$ if there exists a collection of investment decisions $x$ such that $\Pi_{m,i}^{\mathrm{inv}}(x_{i},c)\geq\Pi_{m,i}^{\mathrm{inv}}(x_{i}^{\prime},c)$, for all $x_{i}^{\prime}\in\{0,1\}$ and $i\in\mathcal{I}_{\mathrm{inv}}$, and (1) holds.
@@ III-A Common Model Elements
Consider an operation period (e.g., the representative operation period in Fig. 1) where the expected revenue of investors is shaped by short-term conditions in the distribution-level solar energy market. The market consists of buyers (i.e., electricity consumers), sellers [^3] We acknowledge the possibility that electricity prosumers can be both buyers and sellers at the same time. For the market mechanisms that we will consider in this paper, a prosumer can be modeled as the superposition of a buyer and a seller since buying and selling prices are uniform within each market. This observation may not apply to other markets where buying and selling prices are non-uniform. (i.e., the collection of investors who choose to invest), and a utility company that provides supply when solar generation alone cannot fully satisfy demand. Similar to the investment game, we model the groups of a large number of buyers and sellers as continuous intervals through a non-atomic game formulation. [^4] Non-atomic game formulations [20] differ from non-atomistic market models [21], where some players are large enough to exercise market power.
Define the set of buyers by the interval $\mathcal{I}_{\mathrm{b}}:=[0,1]$. In our analysis, we consider a situation where environmentally conscious electricity consumers have an extra heterogeneous valuation for solar energy over the conventional energy supplied by the utility. For a buyer $i\in\mathcal{I}_{\mathrm{b}}$, we denote their fixed extra valuation for each unit of load served by solar as $v_{i}\in\mathbb{R}_{+}$, referred to as the solar premium. Accordingly, their willingness to pay for a unit of solar-served load is $\pi_{\mathrm{u}}+v_{i}$. Heterogeneity is captured by the distribution of $\{v_{i}\}_{i\in\mathcal{I}_{\mathrm{b}}}$ across buyers, characterized by a function $F_{V}$ defined as $F_{V}(v)=\int_{i\in\mathcal{I}_{\mathrm{b}}}\,\mathbb{1}\{v_{i}\leq v\}\,\,\mathrm{d}i$, where $\mathbb{1}$ is the indicator function. For simplicity, we assume that $F_{V}$ is differentiable and strictly increasing on $0,\overline{v}]$, where $\overline{v}\geq 0$. We denote its derivative by $f_{V}$, supported on $0,\overline{v}]$, and assume $f_{V}(v)$ is bounded away from zero for all $v\in[0,\overline{v}]$. Let the total electric load of all consumers be $L$. We assume $L$ is inelastic, deterministic, and equally distributed across all consumers.